Find whether the following series are convergent or not


i. ∞Σ n=1 (3n-1)/7^n

1
Expert's answer
2022-02-28T15:34:22-0500

Use the Ratio Test


"|\dfrac{a_{n+1}}{a_n}|=\dfrac{\dfrac{3(n+1)-1}{7^{n+1}}}{\dfrac{3n-1}{7^n}}"

"=\dfrac{1}{7}\cdot \dfrac{3n+2}{3n-1}\to\dfrac{1}{7}\ as\ n\to \infin"



Since "\dfrac{1}{7}<1," then the series "\displaystyle\sum_{n=1}^{\infin}\dfrac{3n-1}{7^n}" is convergent by the Ratio Test.